OpenStudy (anonymous):

You toss 5 coins. What is the probability that all 5 show heads? 1/8 1/32 1/16 1/12 1/4

OpenStudy (sidsiddhartha):

each coin has a 1 in 2 chance of landing heads so probability for each one is=1/2 and u have tosses 5 coins so for 5 coins it will be =(1/2)*(1/2)*(1/2)*(1/2)*(1/2)=1/32